A total of $6000 is deposited into two simple interest accounts. One account has an interest rate of 7.5% and the other has a rate of 11.4%. How much should be invested in the 11.4% account to earn a total interest of $528?
Answer by checkley77(12569)

.114*X+.075(6,000-X)=528
.114X+450-.075X=528
.039X=528-450
.039X=78
X=78/.039
X=2,000 IS THE AMOUNT INVESTED @ 11.4%.
6,000-2,000=4,000 IS THE AMOUNT INVESTED @ 7.5%
PROOF:
.114*2,000+.075*4,000=528
228+300=528
528=528